A circuit shows two resistors whose resistance are 0.40ohms and 0.60ohms connected in parallel. They are part of a circuit which carries a current I, and the current in the 0.40ohms resistor is 1.50A

Calculate:
1) The current in the circuit
2) The current in the 0.60ohm resistor
1
Expert's answer
2019-08-26T10:56:55-0400

The voltage, at the resistance of 0.40 Ohm is calculated using Ohm's Law.


"V_{0.40\Omega}=I_{0.40\Omega}*R"


  • Where

Electric resistance "R=0.40\Omega"

Electric current "I_{0.40\Omega}=1.50A"


Numerically evaluating "V_{0.40\Omega}=1.50A*0.40\Omega=0.60V"


As it is a parallel circuit, the voltage across all resistors is the same, that is.


"V_{0.40\Omega}=V_{0.60\Omega}=0.60V"


Now, calculating the current by the resistance of 0.60Ohm using ohm's law


"I_{0.60\Omega}=\frac{V_{0.60\Omega}}{R}"


  • Where

Voltage "V_{0.60\Omega}=0.60V"

Resistance "R=0.60\Omega"


Numerically evaluating "I_{0.60\Omega}=\frac{0.60V}{0.60\Omega}=1.0A"


Finally


Question (a)

The current that passes through the circuit is

"I=I_{0.40\Omega}+I_{0.60\Omega}=1.50A+1.0A=\boxed{2.5A}"


Question (b)

The current passing through the resistance of 0.60Ohm is

"I_{0.60\Omega}=\boxed{1.0A}"